OpenStudy (anonymous):

Sub. Formula to evaluate the integral. posted below (2 problems)

OpenStudy (anonymous):

problem one \[\int\limits_{0}^{\pi} (1 + \cos5t)^2 \sin5t dt\] problem two \[\int\limits_{\pi/6}^{\pi/2} \cot x \csc^3 x dx\]

OpenStudy (sirm3d):

first problem, let \[\large u=1+\cos 5t;\space du=-5\sin 5tdt\] second problem, rewrite the problem as \[\large \int\limits_{\pi/6}^{\pi/2}\csc ^2 x (\csc x \cot x dx)\] and use the substitution \[\large u=\csc x;\space du=-\csc x \cot x dx\] in both cases, after integration, return to the original variable x then sub the limits of integration

OpenStudy (anonymous):

oh alright, thanks a lot